# Selective Classification Under Distribution Shifts

### v1

@@ Abstract
In selective classification (SC), a classifier abstains from making predictions that are likely to be wrong to avoid excessive errors. To deploy imperfect classifiers—imperfect either due to intrinsic statistical noise of data or for robustness issue of the classifier or beyond—in high-stakes scenarios, SC appears to be an attractive and necessary path to follow. Despite decades of research in SC, most previous SC methods still focus on the ideal statistical setting only, i.e., the data distribution at deployment is the same as that of training, although practical data can come from the wild. To bridge this gap, in this paper, we propose an SC framework that takes into account distribution shifts, termed generalized selective classification, that covers label-shifted (or out-of-distribution) and covariate-shifted samples, in addition to typical in-distribution samples, the first of its kind in the SC literature. We focus on non-training-based confidence-score functions for generalized SC on deep learning (DL) classifiers, and propose two novel margin-based score functions. Through extensive analysis and experiments, we show that our proposed score functions are more effective and reliable than the existing ones for generalized SC on a variety of classification tasks and DL classifiers. The code is available at https://github.com/sun-umn/sc_with_distshift.
@@ 1 Introduction
To achieve minimal and controllable levels of classification error so that imperfect and unrobust classifiers can be deployed for high-stakes applications, a promising approach is selective classification (SC): samples that are likely to be misclassified are selected, excluded from prediction, and deferred to human decision makers, so that the classification performance on the remaining samples reaches the desired level (Chow, 1970; Franc et al., 2023; Geifman & El-Yaniv, 2017). For example, by flagging and passing uncertain patient cases that it tends to mistake on to human doctors, an intelligent medical agent can make confident and correct diagnoses for the rest. This “conservative” classification framework not only saves doctors’ efforts, but also avoids liability due to the agent’s mistakes.
@@
It is clear that in practical deployment of classifiers, samples can come from the wild, and hence Type A, Type B and Type C errors can coexist. So, to ensure the reliable deployment of classifiers in high-stakes applications, we must control the three types of errors, jointly. Unfortunately, previous research falls short of a unified treatment of these errors. Classical SC (Chow, 1970) focuses on rejecting samples that cause In-D errors (Type A), whereas the current out-of-distribution (OOD) detection research (Yang et al., 2021; Park et al., 2023) focuses on detecting label-shifted samples (Type B). Although Hendrycks & Gimpel (2016); Granese et al. (2021); Xia & Bouganis (2022); Kim et al. (2023) have advocated the simultaneous detection of samples that cause Type A and Type B errors, their approaches still treat the problem as consisting of two separate tasks, reflected in their separate and independent performance evaluation on OOD detection and SC. Regarding the challenge posed by Type C errors, existing work (Hendrycks & Dietterich, 2018; Croce et al., 2020) focuses primarily on obtaining classifiers that are more robust to covariate shifts, not on rejecting potentially misclassified samples due to covariate shifts—the latter, to the best of our knowledge, has not yet been explicitly considered, not to mention joint rejection together with Type A and Type B errors.
Our contributions
@@
  • We propose and advocate a new SC framework, generalized selective classification, which rejects samples that could cause Type A, Type B and Type C errors jointly, to improve classification performance over the non-rejected samples. With careful review and reasoning, we argue that generalized SC covers and unifies the scope of the existing OOD detection and SC, if the goal is to achieve reliable classification on the selected samples. (Sections 2.3 and 2.4)
@@
  • We propose two confidence scores based on the raw logits (vs. the normalized logits, i.e., softmax responses), inspired by the notion of margins (Section 3.2). Through careful analysis (Section 3.3) and extensive experiments (Section 4), we show that our margin-based confidence scores are more reliable for generalized SC on various dataset-classifier combinations, even under moderate distribution shifts.
@@ 2.2 Prior work in SC
where $\varepsilon\in[0,1]$ is the cost of making a rejection. The optimal selective classifier for this formulation is (Chow, 1970; Franc et al., 2023):
@@
also depend on the posterior probabilities (Pietraszek, 2005; Geifman & El-Yaniv, 2017; Franc et al., 2023; El-Yaniv et al., 2010).
@@
Due to the intractability of true posterior probabilities in practice, many previous methods focus on learning effective confidence-score functions from training data. They require access to training data and learn parametric score functions, often under cost-based/constrained formulations and their variants for the RC tradeoff. This learning problem can be formulated together with (Chow, 1970; Pietraszek, 2005; Grandvalet et al., 2008; El-Yaniv et al., 2010; Cortes et al., 2016; Geifman & El-Yaniv, 2019; Liu et al., 2019; Huang et al., 2022; Gal & Ghahramani, 2016; Lakshminarayanan et al., 2017; Geifman et al., 2018; Maddox et al., 2019; Dusenberry et al., 2020; Lei, 2014; Villmann et al., 2016; Corbière et al., 2019) or separately from training the classifier (Jiang et al., 2018; Fisch et al., 2022; Franc et al., 2023). However, Feng et al. (2023) has recently shown that these training-based scores do not outperform simple non-training-based scores described below.
@@
This family works with any given classifier and does not assume access to the training set. This is particularly attractive when it comes to modern pretrained large DNN models, e.g., CLIP (Radford et al., 2021), Florence (Yuan et al., 2021), and GPTs (Brown et al., 2020), for which obtaining the original training data and performing retraining are prohibitively expensive, if not impossible, to typical users. Algorithm 1 shows a typical use case of SC with non-training-based scores. Different confidence scores have been proposed in the literature. For example, for support vector machines (SVMs), confidence margin (the difference of the top two raw logits) has been used as a confidence score (Fumera & Roli, 2002; Franc et al., 2023); see also Section 3.2. For DNN models, which is our focus, confidence scores are popularly defined over the softmax responses (SRs). Assume that $\mathbf{z}\in\mathbb{R}^{K}$ contains the raw logits (RLs) and $\sigma$ is the softmax activation. The following three confidence-score functions
@@
are popularly used in recent work, e.g., Feng et al. (2023); Granese et al. (2021); Xia & Bouganis (2022). Although simple, $SR_{\text{max}}$ can easily beat existing training-based methods (Feng et al., 2023). On the other hand, these SR-based score functions generally follow the plug-in principle by assuming that SRs approximate posterior probabilities well (Franc et al., 2023). Unfortunately, this assumption often does not hold in practice, and bridging this approximation gap is a major challenge for confidence calibration (Guo et al., 2017; Nixon et al., 2019). However, Zhu et al. (2022) reveals that recent calibration methods may even degrade SC performance.
@@ 2.3 SC under distribution shifts: generalized SC
where $\mathcal{D}_{\mathcal{X},\mathcal{Y}}$ is the original data distribution, $\mathcal{D}^{\prime}_{\mathcal{X},\mathcal{Y}^{\prime}}$ is the shifted distribution—$\mathcal{Y}^{\prime}$ may not be the same as $\mathcal{Y}$ due to potential label shifts. [^1] Here, we assume no outliers—samples that do not follow any specific statistical patterns—during deployment, i.e., they are already detected and removed after separate data preprocessing steps. This allows us to properly define the coverage and selection risk.
@@
Although the existing literature on SC is rich (Zhang et al., 2023), research work that considers SC with potential distribution shifts is very recent and focuses only on label shifts: Xia & Bouganis (2022); Kim et al. (2023) perform In-D SC and OOD (label shift) detection together with a confidence score that combines an SC score and an OOD score, but they still evaluate the performance of In-D SC and OOD detection separately. In contrast, our generalized SC framework considers In-D SC, OOD (label shifts) detection, and covariate shifts, jointly, and we also unify the evaluation protocol (see Section 2.4). Weakly related to our work are Kamath et al. (2020) that studies selective question-answering, which is not a classification problem, under covariate shifts, and Fisch et al. (2022) that performs SC to ensure a good confidence calibration for the selected subset under covariate shifts. Although both consider selective prediction under covariate shifts, their overall goals are different than ours, and their learning-based approaches are also in stark contrast to the non-learning-based approach that we advocate.
@@ 2.4 Evaluation of generalized SC
Since the goal of generalized SC is to identify and exclude misclassified samples, for performance evaluation at a fixed cutoff threshold $\gamma$, it is natural to report the coverage—the portion of samples accepted, and the corresponding selection risk—“accuracy” (taken broadly) on accepted samples. It is clear from Eqs. 1 and 2 that for a given pair of classifier $f$ and confidence-score function $s$, the threshold $\gamma$ can be adjusted to achieve different risk-coverage (RC) tradeoffs. By continuously varying $\lambda$, we can plot a risk-coverage (RC) curve El-Yaniv et al. (2010); Franc et al. (2023) to profile the SC performance of $(f,s)$ throughout the entire coverage range $\phi_{\gamma}\in[0,1]$; see Fig. 1 for an example. Generally, the lower the RC curve, the better the SC performance. To obtain a summarizing metric, it is natural to use the area under the RC curve (AURC) (El-Yaniv et al., 2010; Franc et al., 2023). We note that the RC curve and the AURC are also the most widely used evaluation metrics for classical SC—which is not surprising, as the goal of classical SC aligns with that of generalized SC, although generalized SC also allows distribution shifts.
@@ 3.1 Scale sensitivity of SR-based scores
The above observations are not incidental. To see why the curves change with respect to $\lambda$, note that for a given test set $\left\{\mathbf{x}_{i}\right\}$ and a fixed classifier $f$, the RC curve for any score function $s$ is fully determined by the ordering of $s(\mathbf{x}_{i})$’s (Franc et al., 2023). But this ordering is sensitive to the scale of the raw logits for all three SR-based score functions: $SR_{\text{max}}$, $SR_{\text{doctor}}$, and $SR_{\text{ent}}$. Take $s=SR_{\text{max}}$ as an example and consider any sample $\mathbf{x}$ with its corresponding raw logits $\mathbf{z}$ sorted in descending order (i.e., $z^{(1)}\geq z^{(2)}\geq\cdots$) without loss of generality. Then for any scale factor $\lambda>0$ applied to $\mathbf{z}$, we have the score
@@ 3.2 Our method: margin-based confidence scores
In linear SVMs for binary classification, the classifier takes the form $f(\mathbf{x})=\operatorname{sign}(\mathbf{w}^{\intercal}\mathbf{x}+b)$ and the confidence in classifying a sample $\mathbf{x}$ can be assessed by its distance from the supporting hyperplane (Fumera & Roli, 2002; Franc et al., 2023): |$\absolutevalue{\mathbf{w}^{\intercal}\mathbf{x}+b}/\norm{\mathbf{w}}_{2}$, which is called the geometric margin; see Appendix A for a detailed review. We can extend the idea to $K$-class linear SVMs. Following the popular joint multiclass SVM formulation (Crammer & Singer, 2001), we consider a linear classifier $f(\mathbf{x})=\mathbf{W}^{\intercal}\mathbf{x}+\mathbf{b}$. Here, $\mathbf{W}$ and $\mathbf{b}$ induce $K$ hyperplanes, and we can define the signed distance of any sample $\mathbf{x}$ to the $i$-th yperplanes as: ($\pqty{\mathbf{w}_{i}^{\intercal}\mathbf{x}+b_{i}}/\norm{\mathbf{w}_{i}}$ ($\mathbf{w}_{i}$ denotes the $i$-th column of $\mathbf{W}$ and $b_{i}$ the $i$-th element of $\mathbf{b}$), generalizing the definition for the binary case. However, a single signed distance makes little sense for assessing the classification confidence in multiclass cases, given the typical argmax decision rule—e.g., the largest signed distance can be negative. Instead, comparing the distances to all decision hyperplanes seems more reasonable. Thus, we can consider the following geometric margin as a confidence-score function:
@@
(a-1) RC curves (b-1) $RL_{\text{geo-M}}$ (c-1) $SR_{\text{max}}$ (d-1) Robustness radius
@@
(a-2) RC curves (b-2) Sample visualization (a-3) RC curves (b-3) Sample visualization
Figure 3: Further analysis of the numerical example in Section 3.1. The numbers -$1$, -$2$ and -$3$ in the captions indicate Case 1, 2 and 3, respectively. Here, (a-1), (a-2) and (a-3) plot the RC curves for each case. (b-1), (b-2) and (b-3) visualize the samples (one color per class) and the decision boundaries (dashed blue line), respectively. (b-1) and (c-1) visualize the rejected samples (black crosses) at coverage $0.8$ in Case 1 by $RL_{\text{geo-M}}$ and $SR_{\text{max}}$, respectively. (d-1) shows the histogram of the robustness radius of the selected samples in Case 1 by all score functions.

### v2

@@ Abstract
In selective classification (SC), a classifier abstains from making predictions that are likely to be wrong to avoid excessive errors. To deploy imperfect classifiers—either due to intrinsic statistical noise of data or for robustness issue of the classifier or beyond—in high-stakes scenarios, SC appears to be an attractive and necessary path to follow. Despite decades of research in SC, most previous SC methods still focus on the ideal statistical setting only, i.e., the data distribution at deployment is the same as that of training, although practical data can come from the wild. To bridge this gap, in this paper, we propose an SC framework that takes into account distribution shifts, termed generalized selective classification, that covers label-shifted (or out-of-distribution) and covariate-shifted samples, in addition to typical in-distribution samples, the first of its kind in the SC literature. We focus on non-training-based confidence-score functions for generalized SC on deep learning (DL) classifiers, and propose two novel margin-based score functions. Through extensive analysis and experiments, we show that our proposed score functions are more effective and reliable than the existing ones for generalized SC on a variety of classification tasks and DL classifiers. The code is available at https://github.com/sun-umn/sc_with_distshift.
@@ 1 Introduction
To achieve minimal and controllable levels of classification error so that imperfect and unrobust classifiers can be deployed for high-stakes applications, a promising approach is selective classification (SC): samples that are likely to be misclassified are selected, excluded from prediction, and deferred to human decision makers, so that the classification performance on the remaining samples reaches the desired level (Chow, 1970; Franc et al., 2023a; Geifman & El-Yaniv, 2017). For example, by flagging and passing uncertain patient cases that it tends to mistake on to human doctors, an intelligent medical agent can make confident and correct diagnoses for the rest. This “conservative” classification framework not only saves doctors’ efforts, but also avoids liability due to the agent’s mistakes.
@@
It is clear that in practical deployment of classifiers, samples can come from the wild, and hence Type A, Type B and Type C errors can coexist. In order to ensure the reliable deployment of classifiers in high-stakes applications, we must control the three types of errors, jointly. Unfortunately, previous research falls short of a unified treatment of these errors. Classical SC (Chow, 1970) focuses on rejecting samples that cause In-D errors (Type A), whereas the current out-of-distribution (OOD) detection research (Yang et al., 2021; Park et al., 2023) focuses on detecting label-shifted samples (Type B). Although Hendrycks & Gimpel (2016); Granese et al. (2021); Xia & Bouganis (2022); Kim et al. (2023) have advocated the simultaneous detection of samples that cause Type A and Type B errors, their approaches still treat the problem as consisting of two separate tasks, reflected in their separate and independent performance evaluation on OOD detection and SC. Regarding the challenge posed by Type C errors, existing work (Hendrycks & Dietterich, 2018; Croce et al., 2020) focuses primarily on obtaining classifiers that are more robust to covariate shifts, not on rejecting potentially misclassified samples due to covariate shifts—the latter, to the best of our knowledge, has not yet been explicitly considered, not to mention joint rejection together with Type A and Type B errors.
@@
  • We advocate a new SC framework, generalized selective classification, which rejects samples that could cause Type A, Type B and Type C errors jointly, to improve classification performance over the non-rejected samples. With careful review and reasoning, we argue that generalized SC covers and unifies the scope of the existing OOD detection and SC, if the goal is to achieve reliable classification on the selected samples. (Sections 2.3 and 2.4)
@@
  • We propose two confidence scores based on the raw logits (v.s. the normalized logits, i.e., softmax responses), inspired by the notion of margins (Section 3.2). Through careful analysis (Section 3.3) and extensive experiments (Section 4), we show that our margin-based confidence scores are more reliable for generalized SC on various dataset-classifier combinations, even under moderate distribution shifts.
@@ 2.2 Prior work in SC
where $\varepsilon\in[0,1]$ is the cost of making a rejection. The optimal selective classifier for this formulation is (Chow, 1970; Franc et al., 2023a):
@@
also depend on the posterior probabilities (Pietraszek, 2005; Geifman & El-Yaniv, 2017; Franc et al., 2023a; El-Yaniv et al., 2010).
@@
Due to the intractability of true posterior probabilities in practice, many previous methods focus on learning effective confidence-score functions from training data. They require access to training data and learn parametric score functions, often under cost-based/constrained formulations and their variants for the RC tradeoff. This learning problem can be formulated together with (Chow, 1970; Pietraszek, 2005; Grandvalet et al., 2008; El-Yaniv et al., 2010; Cortes et al., 2016; Geifman & El-Yaniv, 2019; Liu et al., 2019; Huang et al., 2022; Gal & Ghahramani, 2016; Lakshminarayanan et al., 2017; Geifman et al., 2018; Maddox et al., 2019; Dusenberry et al., 2020; Lei, 2014; Villmann et al., 2016; Corbière et al., 2019) or separately from training the classifier (Jiang et al., 2018; Fisch et al., 2022; Franc et al., 2023a). However, Feng et al. (2023) has recently shown that these training-based scores do not outperform simple non-training-based scores described below.
@@
This family works with any given classifier and does not assume access to the training set. This is particularly attractive when it comes to modern pretrained large DNN models, e.g., CLIP (Radford et al., 2021), Florence (Yuan et al., 2021), and GPTs (Brown et al., 2020), for which obtaining the original training data and performing retraining are prohibitively expensive, if not impossible, to typical users. Algorithm 1 shows a typical use case of SC with non-training-based scores. Different confidence scores have been proposed in the literature. For example, for support vector machines (SVMs), confidence margin (the difference of the top two raw logits) has been used as a confidence score (Fumera & Roli, 2002; Franc et al., 2023a); see also Section 3.2. For DNN models, which is our focus, confidence scores are popularly defined over the softmax responses (SRs). Assume that $\mathbf{z}\in\mathbb{R}^{K}$ contains the raw logits (RLs) and $\sigma$ is the softmax activation. The following three confidence-score functions
@@
are popularly used in recent work, e.g., Feng et al. (2023); Granese et al. (2021); Xia & Bouganis (2022). Although simple, $SR_{\text{max}}$ can easily beat existing training-based methods (Feng et al., 2023). On the other hand, these SR-based score functions generally follow the plug-in principle by assuming that SRs approximate posterior probabilities well (Franc et al., 2023a). Unfortunately, this assumption often does not hold in practice, and bridging this approximation gap is a major challenge for confidence calibration (Guo et al., 2017; Nixon et al., 2019). However, Zhu et al. (2022) reveals that recent calibration methods may even degrade SC performance.
@@ 2.3 SC under distribution shifts: generalized SC
where $\mathcal{D}_{\mathcal{X},\mathcal{Y}}$ is the original data distribution, $\mathcal{D}^{\prime}_{\mathcal{X},\mathcal{Y}^{\prime}}$ is the shifted distribution—$\mathcal{Y}^{\prime}$ may not be the same as $\mathcal{Y}$ due to potential label shifts. [^1] We assume no outliers in generalized SC—samples that do not follow any specific statistical patterns—during deployment, i.e., they are already detected and removed after separate data preprocessing steps. This allows us to properly define the coverage and selection risk.
@@
Other related concepts
Besides OOD detection, OOD generalization focuses on correctly classifying In-D and covariate-shifted samples, without considering prediction confidence and selection to improve prediction reliability; open-set recognition (OSR) focuses on correctly classifying In-D samples, as well as flagging label-shifted samples; see Geng et al. (2020) for a comprehensive review. In contrast, generalized SC covers all In-D, label-shifted, and covariate-shifted samples, the widest coverage compared to these related concepts, and targets the most practical and pragmatic metric—classification performance on the selected samples.
@@
Although the existing literature on SC is rich (Zhang et al., 2023), research work that considers SC with potential distribution shifts is very recent and focuses only on label shifts: Xia & Bouganis (2022); Kim et al. (2023) perform In-D SC and OOD (label shift) detection together with a confidence score that combines an SC score and an OOD score, but they still evaluate the performance of In-D SC and OOD detection separately. Müller et al. (2023); Cattelan & Silva empirically show that existing OOD scores are not good enough for SC tasks with covariate/label-shifted samples; Cattelan & Silva proposes ways to refine these scores with the help of additional datasets to optimize performance. Franc et al. (2024) provides theoretical insights on SC with In-D and label-shifted samples. In contrast, we focus on identifying better confidence scores for generalized SC—that covers both In-D and covariate/label-shifted samples and maximizes the utility of the classifier, and unify the evaluation protocol (see Section 2.4).
@@ 2.4 Evaluation of generalized SC
Since the goal of generalized SC is to identify and exclude misclassified samples, for performance evaluation at a fixed cutoff threshold $\gamma$, it is natural to report the coverage—the portion of samples accepted, and the corresponding selection risk—“accuracy” (taken broadly) on accepted samples. It is clear from Eqs. 1 and 2 that for a given pair of classifier $f$ and confidence-score function $s$, the threshold $\gamma$ can be adjusted to achieve different risk-coverage (RC) tradeoffs. By continuously varying $\gamma$, we can plot a risk-coverage (RC) curve El-Yaniv et al. (2010); Franc et al. (2023a) to profile the SC performance of $(f,s)$ throughout the entire coverage range $\phi_{\gamma}\in[0,1]$; see Fig. 1 for an example. Generally, the lower the RC curve, the better the SC performance. To obtain a summarizing metric, it is natural to use the area under the RC curve (AURC) (El-Yaniv et al., 2010; Franc et al., 2023a). We note that the RC curve and the AURC are also the most widely used evaluation metrics for classical SC—which is not surprising, as the goal of classical SC aligns with that of generalized SC, although generalized SC also allows distribution shifts.
@@ 3.1 Scale sensitivity of SR-based scores
The above observations are not incidental. To see why the curves change with respect to $\lambda$, note that for a given test set $\left\{\mathbf{x}_{i}\right\}$ and a fixed classifier $f$, the RC curve for any score function $s$ is fully determined by the ordering of $s(\mathbf{x}_{i})$’s (Franc et al., 2023a). But this ordering is sensitive to the scale of the raw logits for all three SR-based score functions: $SR_{\text{max}}$, $SR_{\text{doctor}}$, and $SR_{\text{ent}}$. Take $s=SR_{\text{max}}$ as an example and consider any sample $\mathbf{x}$ with its corresponding raw logits $\mathbf{z}$ sorted in descending order (i.e., $z^{(1)}\geq z^{(2)}\geq\cdots$) without loss of generality. Then for any scale factor $\lambda>0$ applied to $\mathbf{z}$, we have the score
@@ 3.2 Our method: margin-based confidence scores
In linear SVMs for binary classification, the classifier takes the form $f(\mathbf{x})=\operatorname{sign}(\mathbf{w}^{\intercal}\mathbf{x}+b)$ and the confidence in classifying a sample $\mathbf{x}$ can be assessed by its distance from the supporting hyperplane (Fumera & Roli, 2002; Franc et al., 2023a): |$\absolutevalue{\mathbf{w}^{\intercal}\mathbf{x}+b}/\norm{\mathbf{w}}_{2}$, which is called the geometric margin; see Appendix A for a detailed review. We can extend the idea to $K$-class linear SVMs. Following the popular joint multiclass SVM formulation (Crammer & Singer, 2001), we consider a linear classifier $f(\mathbf{x})=\mathbf{W}^{\intercal}\mathbf{x}+\mathbf{b}$. Here, $\mathbf{W}$ and $\mathbf{b}$ induce $K$ hyperplanes, and we can define the signed distance of any sample $\mathbf{x}$ to the $i$-th hyperplane as: ($\pqty{\mathbf{w}_{i}^{\intercal}\mathbf{x}+b_{i}}/\norm{\mathbf{w}_{i}}$ ($\mathbf{w}_{i}$ denotes the $i$-th column of $\mathbf{W}$ and $b_{i}$ the $i$-th element of $\mathbf{b}$), generalizing the definition for the binary case. However, a single signed distance makes little sense for assessing the classification confidence in multiclass cases, given the typical argmax decision rule—e.g., the largest signed distance can be negative. Instead, comparing the distances to all decision hyperplanes seems more reasonable. Thus, we can consider the following geometric margin as a confidence-score function:
@@
(a-1) (b-1) (c-1) (d-1)
[…]
Case 2
@@
(a-2) (b-2) (c-2) (d-2)
[…]
Case 3
Refer to caption Refer to caption Refer to caption Refer to caption
(a-3) (b-3) (c-3) (d-3)
[…]
Figure 3: Further analysis of the numerical example in Section 3.1. Case 1, Case 2, and Case 3 correspond to the original dataset in Section 3.1, the dataset after small perturbations, and the dataset after substantial perturbations, respectively. Here, (a-)’s are the RC curves achieved by different selection scores; (b-)’s are visualizations of the samples (one color per class), decision boundaries (dashed blue line) and the rejected samples (black crosses) at coverage $0.8$ by $RL_{\text{geo-M}}$; (c-)’s visualize the rejected samples (black crosses) at coverage $0.8$ by $SR_{\text{max}}$; and (d-)’s present the histogram of the robustness radius of the selected samples in by all score functions.